OpenStudy (anonymous):

-4(x+2)>3x+20 solve the write as interval. I need help writing this in interval

OpenStudy (anonymous):

\[-4x-8>3x+20\] \[12>7x\] \[\frac{12}{7}>x\] \[(-\infty, \frac{12}{7})\]

OpenStudy (anonymous):

okay thanks. for some reason i got -4>x

OpenStudy (amistre64):

-8 -20 = 12?

OpenStudy (amistre64):

-4 > x is correct

OpenStudy (anonymous):

it is lol okay thanks

OpenStudy (amistre64):

if we spin that about so that the arrow head is pointing in the lesser direction: x < -4 is an easier view of it

OpenStudy (anonymous):

if i write it in interval would i get -infinite, -4?

OpenStudy (amistre64):

(-inf, -4) would be the interval notation

OpenStudy (anonymous):

thanks

OpenStudy (amistre64):

consider the [ ] to be equal signs connectd by a bar

OpenStudy (amistre64):



OpenStudy (amistre64):



OpenStudy (amistre64):

that way when something is "or equal to" you know what to use an equal sign connected with a bar

OpenStudy (anonymous):

okay

OpenStudy (amistre64):

lol, it just helps me remember the notation :)

OpenStudy (anonymous):

good way to do things

OpenStudy (anonymous):

oops smote by arithmetic again